OpenStudy (alivejeremy):

Bill and Ron build towers using blocks. The expression 4b + 2 represents the number of blocks Bill uses. The expression 8b – 5 represents the number of blocks Ron uses. Complete the expression to represent the total number of blocks Bill and Ron use.

OpenStudy (anonymous):

Hi, are you ready to help me solve it?

OpenStudy (anonymous):

4b+8b-5+2

OpenStudy (alivejeremy):

yea i guess

OpenStudy (anonymous):

I think

OpenStudy (alivejeremy):

thats how it really look

OpenStudy (alivejeremy):

^

OpenStudy (anonymous):

so do 8b+4b and 5+2

OpenStudy (anonymous):

12b - 7

OpenStudy (alivejeremy):

i think the answer is 12b-3

OpenStudy (alivejeremy):

@Michele_Laino

OpenStudy (anonymous):

Your answer is correct! you did it.

OpenStudy (anonymous):

I think it is 12b-7 am I right?

OpenStudy (anonymous):

wait who?

OpenStudy (anonymous):

@alivejeremy is correct.

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

hint: here we have to do this computation: \[\left( {4b + 2} \right) + \left( {8b - 5} \right)\] which can be rewritten as below: \[\begin{gathered} \left( {4b + 2} \right) + \left( {8b - 5} \right) = \hfill \\ \hfill \\ = \left( {4b + 8b} \right) + \left( {2 - 5} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

^_^

OpenStudy (michele_laino):

so, I confirm your answer @alivejeremy

OpenStudy (alivejeremy):

do i get a medal^_^

OpenStudy (anonymous):

Sure, for your correct answer. I suggest you take a look at what @Michele_Laino did above and try to understand it. You will need to the same thing so many times in the future.